solve the differential equation dy/dx(x/1-x^2)y=xy^1/2 , y(0)=1
1
Expert's answer
2020-03-16T13:17:19-0400

"y'\cdot \frac{x}{1-x^2}=xy^\frac{1}{2}, \quad y(0)=1\\
\frac{dy}{y^\frac{1}{2}}=(1-x^2)dx\\
\int\frac{dy}{y^\frac{1}{2}}=\int(1-x^2)dx\\
2y^\frac{1}{2}=x-\frac{x^3}{3}+c\\
y(0)=1\implies y=1, x=0."

Then

"2=0-0+c\implies c=2"

Solution of differential equation is

"2y^\frac{1}{2}=x-\frac{x^3}{3}+2"